OpenStudy (anonymous):

please help- what is the inverse of sq root (x) + 1

OpenStudy (anonymous):

Your original equation:\[y = \sqrt{x+1}\] we swap "x" and "y" to get:\[x = \sqrt{y+1}\]now square both sides to get rid of the "square root"\[x^2 = y +1\]subtract 1 from both sides \[y = x^2 -1\] does that answer your question?

OpenStudy (anonymous):

or is your original equation \[y =\sqrt{x} +1 ?\]

OpenStudy (anonymous):

If my last statement is true, the new inverse would to solve for "x" then do the swap. which follows as so: Subtract one from both sides:\[y -1 = \sqrt{x}\] square both sides:\[(y-1)^2 = x\] swap "x" and "y" to get the final equation: \[y = (x-1)^2\]